# Classifying representations of finite classical groups of Lie type of dimension up to $\ell^{4}$

### v1

@@ 6. Some applications
Let =$\mathcal{R}=\{\rho_{p}\}_{p}$ be a family of $p$-adic orthogonal Galois representations $\rho_{p}:G_{\mathbb{Q}}\rightarrow\operatorname{GO}_{n}(\overline{\mathbb{Q}}_{p})$ as in [14, Theorem 2]. Then, for almost all primes $p$, the image of $\overline{\rho}_{p}^{proj}$ is contained in an almost simple group with socle a finite simple group of Lie type in characteristic $p$.
@@
As observed in [14, Section 5], if we only had a best knowledge of the classification of irreducible representations of finite simple groups of Lie type, we could improve Proposition 6.1. In particular, the main result of [14] (Theorem 1.1), states that for almost all primes $p$ and $n=4\varpi$ (where $\varpi$ is a prime number such that $17\leq\varpi\leq 73$), the image of $\overline{\rho}_{p}^{proj}$ is equal to one of the following groups:
@@
[…]
In the same line of thought, using the strategy in Section 5 of [14] combined with Theorems 4.1, 5.9 and 5.10, we may prove the following improvement of Proposition 6.1 for $n=4\varpi$ with $\varpi$ a prime number such that $79\leq\varpi\leq 787$.
@@
Let =$\mathcal{R}=\{\rho_{p}\}_{p}$ be a compatible system of $n$-dimensional orthogonal Galois representations $\rho_{p}:G_{\mathbb{Q}}\rightarrow\operatorname{GO}_{n}(\overline{\mathbb{Q}}_{p})$ as in [14, Theorem 2]. If $n=4\varpi$, where $\varpi$ is a prime number such that $79\leq\varpi\leq 787$, then for almost all $p$ the image of $\overline{\rho}_{p}^{proj}$ is equal to
@@
[…]
Assume that $G$ is of a Lie type different from $A_{\ell}$ ($\ell\leq 7$), $B_{\ell}$ , $C_{\ell}$, $D_{\ell}$ ($\ell\leq 5$) and $G_{2}$. In order to determine all irreducible representations in defining characteristic for groups of a fixed Lie type, we first need to compute all factorizations of $n$ into factors greater than 1. Such factorizations are $4\varpi=4\cdot\varpi=2\cdot 2\cdot\varpi=2\cdot 2\varpi$. Appealing to Tables 6.6-6.53 of [11] and the assumptions on $G$, there are no irreducible representations of dimension 2 or 4. Thus, there are no tensor decomposable representations of dimension $4\varpi$. Moreover, according to Theorems 4.1, 5.9, 5.10 and Tables 6.6-6.53 of [11], the only irreducible tensor indecomposable self-dual representations of dimension $4\varpi$ are the natural representations of $C_{2\varpi}$ and $D_{2\varpi}$. However, the natural representation of $C_{2\varpi}$ is symplectic then the only irreducible $4\varpi$-dimensional orthogonal representation occurs when $G$ is of type $D_{2\varpi}$. ∎
Then, we have the following result:
@@
For almost all prime $p$ and each $n$ of the form $4\varpi$, where $\varpi$ is a prime number such that $79\leq\varpi\leq 787$, we have that at least one of the following groups:
@@
In the previous result, we include the possible occurrence of (6.3), (6.4), and (6.5), for $2\leq\ell\leq 5$, as Galois groups of $\mathbb{Q}$. In the future, computational tools could be useful to compute all the representations of these groups for dimensions $n\leq 3148$. Moreover, it is worth that the result is new for $\varpi>181$ while for $\varpi\leq 181$ the foregoing result follows from [11].
@@
Finally we also emphasize that if we assume the computations on the web page of Frack Lübek [^1] http://www.math.rwth-aachen.de/ Frank.Luebeck/chev/WMSmall/index.html the Theorem 6.2 works for $100002477\leq\varpi<$ 10000 and for $121553037\leq\varpi<$ 12155 if we include groups of type $C_{6}$. In this case, the result is new for $\varpi>283$ and for $\varpi\leq 283$ follows from Lübek’s computations.
@@ 7. Apendix
Next we shall present the multiplicities and lengths of the dominants weights $\mu\preceq\lambda$, for each highest weight $\lambda$ listed in the table in Theorem 5.10. We may remark that these tables provide all the information to compute the dimensions of $V(\lambda)$ which implies admissibility of the foregoing weights $\lambda$.
@@ 7.1. Type $C_{\ell}$
(7.1) $(\alpha_{i},\alpha_{j})=\begin{cases}\phantom{-}2,&\text{ if }i=j\\ -1,&\text{ if }i=j\pm 1\\ \phantom{-}0,&\text{ otherwise },\end{cases}$
@@
$\lambda_{1}$ $=\alpha_{\ell}+2\alpha_{\ell-1}+\cdots+(\ell-1)\alpha_{2}+\ell(1/2)\alpha_{1}$
@@
$(\alpha_{i},\delta)=\begin{cases}2,&\text{ if }i=1\\ 1,&\text{ otherwise }.\\ \end{cases}$
@@
highest weight dominant weight $\mu\preceq\lambda$ multiplicity of $\mu$ length of $\mathcal{W}$-orbit
$3\lambda_{\ell}$ $3\lambda_{\ell}$ $1$ $2\ell$
$\lambda_{\ell-1}+\lambda_{\ell}$ $1$ $4\ell(\ell-1)$
$\lambda_{\ell-2}$ $1$ $32^{2}\ell(\ell-1)(\ell-2)/$ 3
[…]
$4\lambda_{\ell}$ $4\lambda_{\ell}$ $1$ $2\ell$
$\lambda_{\ell-1}+2\lambda_{\ell}$ $1$ $4\ell(\ell-1)$
$2\lambda_{\ell}$ $1$ $2\ell$
$\lambda_{\ell-2}+\lambda_{\ell}$ $1$ $2^{2}\ell(\ell-1)(\ell-2)$
$\lambda_{\ell-3}$ $1$ $242^{4}\ell(\ell-1)(\ell-2)(\ell-3)/$ 24
[…]
00 $\ell(\ell+1)/2$ $1$
$\lambda_{\ell-2}$ $\lambda_{\ell-2}$ $1$ $62^{3}\ell(\ell-1)(\ell-2)/$ 6
@@ 00 $\ell(\ell+1)/2$ $1$
$\lambda_{\ell-3}$ $\lambda_{\ell-3}$ $1$ $242^{4}\ell(\ell-1)(\ell-2)(\ell-3)/$ 24
$\lambda_{\ell-1}$ $\ell-3$ $62\ell(\ell-1)/$ 6
00 $\ell(\ell-3)/2$ $1$
$\lambda_{\ell-1}+\lambda_{\ell}$ $\lambda_{\ell-1}+\lambda_{\ell}$ $1$ $2^{2}\ell(\ell-1)$
$\lambda_{\ell-2}$ $2$ $32^{2}\ell(\ell-1)(\ell-2)/$ 3
$\lambda_{\ell}$ $2(\ell-1)$ $2\ell$
@@ 7.2. Type $D_{\ell}$
(7.2) $(\alpha_{i},\alpha_{j})=\begin{cases}\phantom{-}2,&\text{ if }i=j\\ -1,&\text{ if }i=j\pm 1\\ \phantom{-}0,&\text{ otherwise },\\ \end{cases}$
@@
$\lambda_{1}$ $=(1/2)(\alpha_{\ell}+2\alpha_{\ell-1}+\cdots+(\ell-2)\alpha_{3}+(1/2)(\ell-2)\alpha_{2}+(1/2)\ell\alpha_{1})$
@@
$(\alpha_{i},\delta)=\begin{cases}2&\text{ if }i=1,2\\ 1&\text{ otherwise }.\\ \end{cases}$
@@
highest weight dominant weight $\mu\preceq\lambda$ multiplicity of $\mu$ length of $\mathcal{W}$-orbit
$3\lambda_{\ell}$ $3\lambda_{\ell}$ $1$ $2\ell$
@@
$4\lambda_{\ell}$ $4\lambda_{\ell}$ $1$ $2\ell$
$\lambda_{\ell-1}+2\lambda_{\ell}$ $1$ $4\ell(\ell-1)$
$2\lambda_{\ell-1}$ $1$ $2\ell(\ell-1)$
$\lambda_{\ell-2}+\lambda_{\ell}$ $1$ $2^{2}\ell(\ell-1)(\ell-2)$
$\lambda_{\ell-3}$ $1$ $242^{2}\ell(\ell-1)(\ell-2)(\ell-3)/$ 24
[…]

### v3

@@ 5.1. Type $C_{\ell}$ and $D_{\ell}$
Let $G$ of type $C_{\ell}$ over $\overline{\mathbb{F}}_{p}$ with $p>2$. Then,
  1. (1)
    if $p\nmid(\ell-1)!$ then
    $\quad\dim L(\lambda_{\ell-2})=\binom{2\ell}{3}-\binom{2\ell}{1}$
  2. (2)
    if $p\nmid\ell!$ then
    $\quad\dim L(\lambda_{\ell-3})=\binom{2\ell}{4}-\binom{2\ell}{2}$,
  3. (3)
    if $p\neq 3$ then
    $\dim L(3\lambda_{\ell})=\binom{2\ell+2}{3}\qquad\mbox{ and }\qquad\dim L(4\lambda_{\ell})=\binom{2\ell+3}{4}$.
Proof.
For (1) and (2) we may argue similarly as in the proof of [24, Theorem 5.1]. First, for $\lambda_{\ell-2}$, we have the decomposition
$\Lambda^{3}(L(A_{\ell-1}))\oplus(\Lambda^{2}(L(A_{\ell-1}))\otimes L(A_{\ell-1})^{*})\oplus(L(A_{\ell-1})\otimes\Lambda^{2}(L(A_{\ell-1})^{*})\oplus\Lambda^{3}(L(A_{\ell-1})^{*})$
However, in this case we can only ensure that $\Lambda^{3}(L(C_{\ell}))$ contains an irreducible constituent of dimension $\binom{2(\ell-1)}{3}-\binom{2(\ell-1)}{1}$, by applying induction, when $p\nmid((\ell-1)-1)!$. Then by using the irreducibility of $\Lambda^{2}(V(A_{\ell-1}))\otimes V(A_{\ell-1})^{*}$ and $V(A_{\ell-1})\otimes\Lambda^{2}(V(A_{\ell-1})^{*})$ for $p\nmid\ell-1$, we obtain our result. Now, for $\lambda_{\ell-3}$, by using a similar decomposition as above, we may compute the dimension of $\Lambda^{4}(L(C_{\ell}))$, where the condition on $p$ comes from the induction on $\ell$ and the terms $\Lambda^{2}(L(A_{\ell-1}))\otimes\Lambda^{2}(L(A_{\ell-1})^{*})$ and $\Lambda^{3}(L(A_{\ell-1}))\otimes L(A_{\ell-1})^{*}$.
For (3) we can also use a similar argument as above, however it follows more easily from [27, Proposition 4.2.2.h]. ∎
For type $D_{\ell}$, in addition to Lübeck’s formulas [24, Theorem 5.1], Burness, Ghandour, Testerman [9, Lemma 2.3.2], Cavallin [13, Table 9] and Martínez [26, Table 4] have obtained the following formulas
$\dim L(\lambda_{1})=L(\lambda_{2})=2^{\ell-1},\quad\dim L(3\lambda_{\ell})=\binom{2\ell+2}{3}-\binom{2\ell}{1}+\epsilon(\ell+1)(2\ell),$
and
$\dim L(\lambda_{\ell-i})=\binom{2\ell}{i+1}\;\mbox{for}\;p\neq 2\;\mbox{and}\;i=2,3.$
Moreover, if $p\neq 2,3$ and $p\nmid(\ell+2)!$, we can obtain the formula
$\quad\dim L(4\lambda_{\ell})=\binom{2\ell+3}{4}-\binom{2\ell+1}{2}$.
as in the previous proposition. The foregoing formulas, together with [24, Theorem 5.1], allow us to compute the explicit dimension of $L(\lambda)$ (at least for sufficiently large primes $p$, explicitly $p>\ell+2$) for all admissible weights in Theorem 4.7 except for $\lambda=\lambda_{\ell-1}+\lambda_{\ell}$. However, it follows from [26, §4] that
(5.1) $\dim L(\lambda)=\binom{2\ell+2}{3}\quad\mbox{for}\quad p\neq 2,3\mbox{ and }p\nmid(2\ell-1)(2\ell+1)$
in both cases. Note that, this formula agrees with Lübeck’s computations for small $\ell$ [24, §6.32-6.48].
For completeness, the dimension of $L(\lambda)$ for type $B_{\ell}$ may be found in [24, §5] for the weights $0,\;\lambda_{\ell},\;2\lambda_{\ell},\;\lambda_{\ell-1}$; in [13, Table 8] for $\lambda=\lambda_{\ell-a+1}$ with $a=3,4,\ell$; and in [26] for $3\lambda_{\ell}$ and $\lambda_{\ell-1}+\lambda_{\ell}$. For the remaining admissible weight $4\lambda_{\ell}$ we believe that:
(5.2) $\dim L(4\lambda_{\ell})=2\ell(2\ell+2)(2\ell+1)(2\ell+7)/24-\epsilon(2\ell+5)\ell(2\ell+3)-\epsilon(2\ell+3)$
The technique used in [24] does not work in this case. Indeed, it seems to be appropriate to attempt a reverse approach, i.e., embed our group into a group of a suitable type and use the information therein.
5.2. Automorphic Galois representations
[…]
On the other hand, let $\rho_{p}:G_{\mathbb{Q}}\rightarrow\operatorname{GO}_{n}(\overline{\mathbb{Q}}_{p})\subseteq\operatorname{GL}_{n}(\overline{\mathbb{Q}}_{p})$ be a $p$-adic orthogonal Galois representation. As we mentioned in the Introduction, examples of this kind of Galois representations can be obtained from the so-called regular algebraic, polarizable, cuspidal automorphic representations $\pi$ of $\operatorname{GL}_{n}(\mathbb{A}_{\mathbb{Q}})$, via the global Langlands correspondence [4, §2.1]. In that case, the large image conjecture for automorphic Galois representations predicts that, for almost all primes $p$, the image of the projectivizations $\rho^{proj}_{p}:G_{\mathbb{Q}}\rightarrow{\rm PGO}_{n}(\overline{\mathbb{F}}_{p})$ of the residual representations of $\rho_{p}:G_{\mathbb{Q}}\rightarrow\operatorname{GO}_{n}(\overline{\mathbb{Q}}_{p})$ should contain $\operatorname{P\Omega}^{\pm}_{n}(\mathbb{F}_{p^{s}})$ for some positive integer $s$. As we also pointed out at that moment, this conjecture is known in very few cases. However, if we impose certain conditions on the local components of $\pi$, we can at least bound the possible images of $\overline{\rho}_{p}$. For example, in [36, Proposition 3.4], the second author proved the following criterion concerning the image of certain families of orthogonal Galois representations.
Proposition 5.2.
@@ 5.2. Automorphic Galois representations
Let $\mathcal{R}=\{\rho_{p}\}_{p}$ be a family of $p$-adic orthogonal Galois representations $\rho_{p}:G_{\mathbb{Q}}\rightarrow\operatorname{GO}_{n}(\overline{\mathbb{Q}}_{p})$ as in [36, Theorem 2]. Then, for almost all primes $p$ (i.e. all but finitely many), the image of $\overline{\rho}_{p}^{proj}$ is contained in an almost simple group with socle a finite simple group of Lie type in characteristic $p$.
@@
Then, a best knowledge of the classification of irreducible representations of finite simple groups of Lie type can help us to improve Proposition 5.2. In particular, the main result of [36] (Theorem 1.1), states that for almost all primes $p$ and $n=4\varpi$ (where $\varpi$ is a prime number such that $17\leq\varpi\leq 73$), the image of $\overline{\rho}_{p}^{proj}$ is equal to one of the following groups:
@@
for some positive integer $s$. The idea of the proof is to look at the classification of projective irreducible representations of finite simple groups of Lie type PGG of dimension $n\leq 292$, given in [24], and prove that there are not irreducible representations of dimension $n$ other than those coming from a finite simple group of type $D_{2\varpi}$ for $p$ sufficiently large and $n=4\varpi$ (see [36, §2] and [35] for the precise conditions imposed on $p$ and $n$). This improving had an interesting consequence concerning to the inverse Galois problem. Explicitly, [36, Corollary 1.2] states that at least one of the groups in (5.3) can be realized as a Galois group over $\mathbb{Q}$ for infinitely many primes $p$ and infinitely many positive integers $s$.
In the same line of thought, using the strategy in Section 5 of [36] combined with Theorems 3.4, 4.6 and 4.7, we may prove the following conditional improvement of Proposition 5.2 for $n=4\varpi$ with $\varpi$ a prime number such that $79\leq\varpi\leq 787$.
@@
Let $\varpi$ be a prime number such that $79\leq\varpi\leq 787$, $n=4\varpi$, and $\mathcal{R}=\{\rho_{p}\}_{p}$ be a compatible system of $n$-dimensional orthogonal Galois representations $\rho_{p}:G_{\mathbb{Q}}\rightarrow\operatorname{GO}_{n}(\overline{\mathbb{Q}}_{p})$ as in [36, Theorem 2]. Assuming that the formula (5.2) is true, we have that for almost all $p$ the image of $\overline{\rho}_{p}^{proj}$ is equal to
@@
The proof closely follows the ideas of [36, Theorem 1.1]. Nevertheless, we will write a brief sketch of the proof for completeness. As we pointed out above, projective representations of PGG in characteristic $p$ are the same as linear representations of $G^{F}$ in characteristic $p$, which can be constructed by restricting highest weight representations of $G$ to $G^{F}$. Thus, according to the discussion in [36, §5] preceding the proof of Theorem 1.1 (which works for any dimension), it suffices to prove that the only irreducible $4\varpi$-dimensional representations $\overline{\rho}_{p}:G^{F}\rightarrow\Omega_{4\varpi}^{\pm}(\mathbb{F}_{q^{\epsilon}})$ that could occur, appear when $G$ is of type $A_{\ell}$ ($\ell\leq 7$), $B_{\ell}$, $C_{\ell}$, $D_{\ell}$ ($\ell\leq 5$), $G_{2}$ or $D_{\ell}$.
@@
Assume that $G$ is of a Lie type different from $A_{\ell}$ ($\ell\leq 7$), $B_{\ell}$ , $C_{\ell}$, $D_{\ell}$ ($\ell\leq 5$) and $G_{2}$. In order to determine all irreducible representations in defining characteristic for groups of a fixed Lie type, we first need to compute all factorizations of $n$ into factors greater than 1. Such factorizations are $4\varpi=4\cdot\varpi=2\cdot 2\cdot\varpi=2\cdot 2\varpi$. Appealing to Tables 6.6-6.53 of [24] and the assumptions on $G$, there are no irreducible representations of dimension 2 or 4. Thus, there are no tensor decomposable representations of dimension $4\varpi$. Moreover, according to Theorems 3.4, 4.6 and 4.7; the dimension formulas in the previous section (including (5.2)); and Tables 6.6-6.53 of [24], the only irreducible tensor indecomposable self-dual representations of dimension $4\varpi$ are the natural representations of $C_{2\varpi}$ and $D_{2\varpi}$. However, the natural representation of $C_{2\varpi}$ is symplectic then the only irreducible $4\varpi$-dimensional orthogonal representation occurs when $G$ is of type $D_{2\varpi}$. ∎
Then, we will have the following interesting consequence concerning to the inverse Galois problem:
Once we prove that the formula (5.2) is true for $p>2\ell+5$, an interesting consequence concerning to the inverse Galois problem can be deduced as follows.
@@
Assuming that the formula (5.2) is true for $p>2\ell+5$, we have that, for almost all prime $p$ and each $n$ of the form $4\varpi$ (where $\varpi$ is a prime number such that $79\leq\varpi\leq 787$), at least one of the following groups:
@@
In the previous result, we include the possible occurrence of (5.5), (5.6), and (5.7), for $2\leq\ell\leq 7$, as Galois groups of $\mathbb{Q}$. In the future, computational tools could be useful to compute all the representations of these groups for dimensions $n\leq 3148$ and obtain the exact analogue of [36, Corollary 1.2]. It is worth pointing out that such a result will be new for $\varpi>181$ while for $\varpi\leq 181$ the foregoing result follows from [24].
@@
Finally we would like also to emphasize that if we assume the computations on the web page of Frack Lübek [^1] http://www.math.rwth-aachen.de/ Frank.Luebeck/chev/WMSmall/index.html the Theorem 5.3 also works for $2477\leq\varpi<10000$ and for $3037\leq\varpi<12155$ if we include groups of type $C_{6}$. In this case, the result will be new for $\varpi>283$ and for $\varpi\leq 283$ follows from Lübek’s computations.
Remark 5.7.
Besides this motivating application to Number Theory, there are many other interesting applications to Algebra itself. See, for example [10, 17, 21, 22, 29, 30]
@@ 6. Apendix
Next we shall present the multiplicities and lengths of the dominants weights $\mu\preceq\lambda$, for each highest weight $\lambda$ listed in the table in Theorem 4.7. We may remark that these tables provide all the information to bound the dimensions of $V(\lambda)$ (and whence $L(\lambda)$) which implies admissibility of the foregoing weights $\lambda$.
@@ 6.1. Type $C_{\ell}$
(6.1) $(\alpha_{i},\alpha_{j})=\begin{cases}\phantom{-}2,&\text{ if }i=j\\ -1,&\text{ if }i=j\pm 1\\ \phantom{-}0,&\text{ otherwise }\end{cases}$
@@
$\lambda_{1}$ $=\alpha_{\ell}+2\alpha_{\ell-1}+\cdots+(\ell-1)\alpha_{2}+\ell(1/2)\alpha_{1},$
@@
[…]
@@
highest weight dominant weight $\mu\preceq\lambda$   bound of $m_{L(\lambda)}(\mu)$ length of $\mathcal{W}$-orbit
$3\lambda_{\ell}$ $3\lambda_{\ell}$ 11 $2\ell$
$\lambda_{\ell-1}+\lambda_{\ell}$ 11 $4\ell(\ell-1)$
$\lambda_{\ell-2}$ 11 $2^{2}\ell(\ell-1)(\ell-2)/3$
$\lambda_{\ell}$ $\ell$ $2\ell$
$4\lambda_{\ell}$ $4\lambda_{\ell}$ 11 $2\ell$
$\lambda_{\ell-1}+2\lambda_{\ell}$ 11 $4\ell(\ell-1)$
$2\lambda_{\ell}$ 11 $2\ell$
$\lambda_{\ell-2}+\lambda_{\ell}$ 11 $2^{2}\ell(\ell-1)(\ell-2)$
$\lambda_{\ell-3}$ 11 $2^{4}\ell(\ell-1)(\ell-2)(\ell-3)/24$
$\lambda_{\ell} \ell 22\ell$
$\lambda_{\ell-1}$ $\ell$ $2\ell(\ell-1)$
0 $\ell(\ell+1)/2$ 11
$\lambda_{\ell-2}$ $\lambda_{\ell-2}$ 11 $2^{3}\ell(\ell-1)(\ell-2)/6$
@@ 0 $\ell(\ell+1)/2$ 11
$\lambda_{\ell-3}$ $\lambda_{\ell-3}$ 11 $2^{4}\ell(\ell-1)(\ell-2)(\ell-3)/24$
$\lambda_{\ell-1}$ $\ell-3$ $2\ell(\ell-1)/6$
0 $\ell(\ell-3)/2$ 11
$\lambda_{\ell-1}+\lambda_{\ell}$ $\lambda_{\ell-1}+\lambda_{\ell}$ 11 $2^{2}\ell(\ell-1)$
$\lambda_{\ell-2}$ 22 $2^{2}\ell(\ell-1)(\ell-2)/3$
$\lambda_{\ell}$ $(\ell-1) 22\ell$
@@ 6.2. Type $D_{\ell}$
[…]
@@
$\lambda_{1}$ $=(1/2)(\alpha_{\ell}+2\alpha_{\ell-1}+\cdots+(\ell-2)\alpha_{3}+(1/2)(\ell-2)\alpha_{2}+(1/2)\ell\alpha_{1}),$
@@
[…]
@@
highest weight dominant weight $\mu\preceq\lambda$   a bound of $m_{L(\lambda)}(\mu)$ length of $\mathcal{W}$-orbit
$3\lambda_{\ell}$ $3\lambda_{\ell}$ 11 $2\ell$
@@
$4\lambda_{\ell}$ $4\lambda_{\ell}$ 11 $2\ell$
$\lambda_{\ell-1}+2\lambda_{\ell}$ 11 $4\ell(\ell-1)$
$2\lambda_{\ell-1}$ 11 $2\ell(\ell-1)$
$\lambda_{\ell-2}+\lambda_{\ell}$ 11 $2^{2}\ell(\ell-1)(\ell-2)$
$\lambda_{\ell-3}$ 11 $2^{2}\ell(\ell-1)(\ell-2)(\ell-3)/24$
$\lambda_{\ell} \ell 22\ell$
$\lambda_{\ell-1}$ $\ell$ $2\ell(\ell-1)$